# ]https://zf-wei.github.io

### v2

@@ II.4 Community similarity
Network Generation, Sec. II.1Perturbation, Sec. II.2Data Clustering, Sec. II.3Similarity, Sec. II.4
@@
III Results and discussion
In this section, we present the results of our experiments. Recall that Sec. II.3 details the embedding methods we use here. In the following subsections, we use $\mathbf{A}$ to denote the adjacency matrix of a graph. We report results based on the family of graph embedding methods (i.e., matrix-based family in Sec. III.1 and random walk family in Sec. III.2).
In the subsequent line figures, unless otherwise stated (in Figs. 16, 24, 25 and 26), subfigures in the top row, i.e., (a) and (b), correspond to random node selection. More specifically, subfigure (a) corresponds to experimental results on small networks with $1,000$ nodes and subfigure (b) depicts experimental results on large networks with $10,000$ nodes. Similarly, subfigures in the bottom row, i.e., (c) and (d), correspond to targeted node selection, for small and large networks respectively.In the body of this paper, we present line figures using $16$-dimensional embedding while line figures with $32$-dimensional embedding are shown in Appendix A. We focus only in these two dimensions as smaller embedding dimensions seem to be enough to provide an accurate low-dimensional vector representation of networks [15, 16]. Recall that a higher ECS value indicates a greater similarity between network community partitions obtained from the LFR benchmark graphs and the embeddings. Thus, when analyzing line figures, higher ECS values and therefore higher positions of line curves indicate a more robust graph embedding method for community detection.
In addition, to facilitate a thorough comparison of embedding methods, heatmaps are employed to offer a more comprehensive assessment on the joint effect of network size and embedding dimension in community robustness for a fixed $\mu$ in absence of perturbations. Specifically, for different community partition strengths (i.e., strong with $\mu=0.01$ and weak with $\mu=0.5$), $50$ independent instances of LFR networks are generated, each with a distinct network size. The identified community structures, with different embedding dimensions, are then compared with the ground-truth community partition using the ECS similarity score. The mean similarity score derived from $50$ copies of LFR networks is utilized to populate each cell in the heatmap. When producing heatmaps, we notice that the standard deviations are negligible with respect to the mean; so, we do not show those results. Note that no edge removal is involved in drawing heatmaps.
Table 2 shows the experimental parameters employed by the embedding methods, excluding embedding dimensions, $d$. Notably, LE and LLE methods only require the embedding dimension $d$. For the remaining parameters, we utilized the default parameters as specified in Tandon et al. (2021) [14]; they show that default values of parameters for the embedding methods lead to comparable performance as traditional community detection methods, and that the optimized parameters do not exhibit a significant improvement over default parameters.
@@ III Results and discussion
Random Walk LINE batch size: $8192$, epochs: $100$
@@
In Sec. III.3, we conduct cross-comparisons between data clustering methods using heatmaps, to compare how similar are the community detection outcomes based on various embedding techniques. We observe that LE and LLE, as two methods in the family of matrix-based methods, yield similar data clustering results; similarly, DeepWalk and node2vec, as two methods in the family of random walk-based approaches, also exhibit similar results.
Finally, Sec. III.4 presents experimental results to evaluate the robustness of our community detection methods on two real-world networks with labeled communities: the email-EU-core network [63] and the AS network [7]. These networks exhibit similar characteristics to synthetic counterparts, specifically power-law degree and community size distributions as observed in LFR benchmark graphs. Our experimental methodology and insights for real-world networks mirror findings for synthetic counterparts.
III.1 Matrix factorization methods
@@ III.1.1 Laplacian eigenmap (LE)
$\mathbf{D}^{-1}(\mathbf{D}-\mathbf{A})$
In Fig. 2, the impact of edge removal based on different node selection strategies is demonstrated. The effects of the initial community partition’s strength (with stronger partitions experiencing less perturbation impact, i.e., lower $\mu$ values) and the type of perturbation (targeted node selection having a significant influence on clusters) are observed. Specifically, for networks with strong community partition (i.e., $\mu=0.01$ or $\mu=0.1$), we notice that embedding dimension tends to have a stronger effect on the decay rate of similarity, despite node selection strategy. For instance, in Fig. 2(a), with an embedding dimension of $16$ under random node selection, the curve corresponding to $\mu=0.1$ remains consistently stable around an ECS value of approximately $1.0$ until $60\%$ of nodes are selected. Subsequently, between $60\%$ and $85\%$ of nodes selected, the curve gradually and steadily decreases from $1.0$ to $0.84$ (i.e., $16\%$ decrease). In contrast, depicted in Fig. 17(a) (see Appendix A) using $32$-dimensional embedding, the curve representing $\mu=0.1$ remains stable around $1.0$ only until $25\%$ of node are selected. Then, between $25\%$ and $85\%$ of nodes selected, the curve decreases more sharply from $1.0$ to $0.53$ (i.e., $47\%$ decrease).
Fig. 3 helps us to better understand the effect of embedding dimension and network size. We notice that for strongly clustered networks (i.e., $\mu=0.01$), a lower embedding dimension produces higher ECS. In contrast, for weakly clustered networks (i.e., $\mu=0.5$), lower embedding dimensions produce higher ECS for networks with small sizes and higher embedding dimensions result in higher ECS for networks with larger sizes. In both scenarios represented by Fig. 3(a) and Fig. 3(b), it is observed that, for a given fixed embedding dimension, the data clustering method employing LE yields higher ECS scores for larger networks.
Figure 2: Mean ECS using LE method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. Random node selection is used in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) are for networks with $1,000$ nodes, and (b) and (d) are for networks with $10,000$ nodes.
Figure 3: Heatmaps with LE embedding method. For given embedding dimension and LFR network with given size, the ECS similarity score between ground-truth community structure and detected community structure is recorded.
III.1.2 Locally linear embedding (LLE)
@@
Figure 4: Mean ECS using LLE method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. (a) and (b) apply random node selection, whereas (c) and (d) use targeted selection. Networks in (a) and (c) have $1,000$ nodes, and those in (b) and (d) have $10,000$ nodes.
The insights drawn from Figs. 4, 18 and 5 for LLE align closely with the insights derived for LE. In fact, the curves in Fig. 4 are slightly higher in position than those in Fig. 2, indicating that LLE outperforms LE. For example, the data points on the green curve (corresponding to $\mu=0.4$) from LLE in Fig. 4(a) have higher ECS scores over LE in Fig. 2(a) for a given percentage of selected nodes.
Figure 5: Heatmaps with LLE embedding method. For given embedding dimension and LFR network with given size, the ECS similarity score between ground-truth community structure and detected community structure is recorded.
III.1.3 Higher-order preserving embedding (HOPE)
@@
Our experimental results on the impact of edge removal, generated through the HOPE method with embedding dimension $16$, are presented in Fig. 6. For the HOPE embedding method, our results suggest that using a 32-dimensional embedding (i.e., Fig. 19 in Appendix A) brings higher ECS scores than a 16-dimensional one (i.e., Fig. 6). For example, the data points on the green curve (corresponding to $\mu=0.4$) in Fig. 19(a) have higher ECS scores than those data points on the green curve (corresponding to $\mu=0.4$) in Fig. 6(a) at respective percentage of selected nodes. We can also see that HOPE has higher ECS scores with smaller graphs. Analysis of heatmaps in Fig. 7 reveals that the HOPE method tends to preserve effectively higher ECS scores for smaller networks and higher embedding dimensions.
Figure 6: Mean ECS using HOPE method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. While (a) and (b) use random node selection, (c) and (d) rely on targeted selection. Networks in (a) and (c) contain $1,000$ nodes, while those in (b) and (d) consist of $10,000$ nodes.
Figure 7: Heatmaps with HOPE embedding method. For given embedding dimension and LFR network with given size, the ECS similarity score between ground-truth community structure and detected community structure is recorded.
III.1.4 Modularized non-negative matrix factorization (M-NMF)
@@
Our experimental results on the impact of edge removal, generated through the M-NMF embedding method with dimension $16$, are reported in Fig. 8. From these figures, we can see that M-NMF produces higher ECS scores with smaller graphs. Our experimental results using $32$-embedding are collected in Fig. 20 (see Appendix A). In Appendix B, we compared these two embedding dimensions for M-NMF. Reviewing heatmaps in Fig. 9, M-NMF produces higher ECS scores for strongly clustered networks, but a clear pattern is not discernible. Notably, from Fig. 9(b), M-NMF produces lower ECS scores in cases where $\mu=0.5$. Recall that M-NMF tries to preserve the community structure of networks and therefore is strongly affected by perturbations when the initial community structure is weak.
Figure 8: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. (a) and (b) implement random node selection, while (c) and (d) apply targeted selection; (a) and (c) correspond to networks with $1,000$ nodes, and (b) and (d) to networks with $10,000$ nodes.
Figure 9: Heatmaps with M-NMF embedding method. For given embedding dimension and LFR network with given size, the ECS similarity score between ground-truth community structure and detected community structure is recorded.
@@
Larger Network (larger $N$) LE with a higher embedding dimension LE with a lower embedding dimension
III.1.5 Summary of matrix factorization methods
We derive the following insights for matrix factorization methods. In a nutshell, we observe that ECS scores consistently decrease as we progressively select more nodes and delete their adjacent edges and the networks become sparser. This aligns with our intuition that the community structure is more significantly disrupted when a greater number of edges are removed. From previous figures, it is evident that, in general, for smaller values of $\mu$, the ECS similarity score tends to remain higher for varying degrees of perturbations. This suggests that the graph embedding methods tend to produce higher ECS scores when the original LFR benchmark graph exhibits a stronger partition. Recall that $\mu$ signifies the fraction of nodes sharing edges across different communities. Thus, a lower $\mu$ value implies a higher ratio of internal to external edges, usually resulting in a more pronounced partition in the LFR benchmark graph. Notably, when utilizing targeted node selection, the ECS exhibits a more pronounced and rapid decline compared to random selection. For example, for LFR network with $1,000$ nodes and $\mu=0.5$, when perturbing the network at $50\%$ and using LE with $16$-dimensional embedding, we notice that targeted node selection decreases ECS from $0.67$ to $0.14$ (a decrease of $79.1\%$, see Fig. 2(c)), but ECS decreases from $0.67$ to $0.35$ for random node selection (a decrease of $47.8\%$, see Fig. 2(a)). This observation is comprehensible, as targeted node selection based on betweenness centrality tends to dismantle the network’s community structure more swiftly.
For smaller real-world graphs containing a few thousand vertices, to obtain more robust partitions, the LLE method is suggested with a modest embedding dimension, typically on the order of tens, such as 16 dimensions; Alternatively, when the community partition is stronger, the HOPE embedding method with a higher embedding dimension produces comparably robust partitions. In the case of larger real-world graphs with a weaker community partition, to obtain more robust community partitions, it is suggested to use the LE embedding method with a moderately higher embedding dimension, such as $32$ or $64$ dimensions as utilized in our experiments (see Fig. 17 in Appendix A). Alternatively, LLE with a higher embedding dimension produces partitions that are similarly robust but marginally inferior. When dealing with larger real-world graphs characterized by a stronger community structure, the LLE or LE embedding method with a lower embedding dimension produces more robust community partitions. In summary, LLE is the most robust method among matrix factorization methods. Table 3 summarizes our findings for matrix factorization methods.
III.2 Random walk methods
Random walk methods learn network embeddings of graph nodes by modeling a stream of short random walks. We chose DeepWalk [31], LINE [32], and node2vec [33], as widely used methods of this family.
III.2.1 DeepWalk
@@
Our experimental results on the impact of edge removal, generated through the DeepWalk embedding method with dimension $16$, are reported in Fig. 10. Our experimental results using DeepWalk embedding method with dimension $32$ are collected in Fig. 21. We can observe that networks with stronger initial partitions (i.e., lower $\mu$ values) experience less perturbation impact and targeted node selection has a significant influence on clusters. Examining heatmaps in Fig. 11(a), it is evident that for strongly clustered networks, the DeepWalk embedding method produces higher ECS scores for smaller network sizes and lower embedding dimensions.
Figure 10: Mean ECS using DeepWalk method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings.Random selection is employed in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) represent networks with $1,000$ nodes, and (b) and (d) represent those with $10,000$ nodes.
Figure 11: Heatmaps with DeepWalk embedding method. For given embedding dimension and LFR network with given size, the ECS similarity score between ground-truth community structure and detected community structure is recorded.
III.2.2 Large-scale information network embedding (LINE)
@@
where $E$ is the edge set and $\mathbf{x}_{i}$ denotes the vector indicating the position of the point representing node $i$ in the embedding.
The optimization process of the LINE method utilizes stochastic gradient descent, which is enhanced by an edge-sampling treatment, as detailed in [32]. Specifically, the utilization of the LINE embedding method necessitates the presence of GPU for computation. In our study, the machine learning model was trained using a batch size of $8192$ over $100$ epochs, facilitating iterative updates to the embedding vector across the entire training dataset.
Our experimental results generated through the LINE embedding method with dimension $16$ are presented in Fig. 12. Our experimental results using LINE embedding method with dimension $32$ are collected in Fig. 22 (see Appendix A). The LINE embedding method produces higher ECS scores when applied to larger networks. Our analysis of networks with 1,000 nodes in the first column of Fig. 12 reveals that curves exhibit a steep decline followed by a subsequent increase, but we notice that increasing the embedding dimensions in LINE is expected to mitigate the sharpness of this decline and increase; see Appendix B for details. Owing to the substantial computational demands of the LINE embedding method, we omit the heatmap analysis specifically for LINE.
Based on the aforementioned observation, LINE becomes a viable option for network clustering when dealing with relatively larger networks and the computational resources, particularly GPUs, are available.
Figure 12: Mean ECS using LINE method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. (a) and (b) follow random node selection, whereas (c) and (d) adopt targeted node selection. Networks with $1,000$ nodes correspond to (a) and (c), and those with $10,000$ nodes correspond to (b) and (d).
III.2.3 node2vec
@@
Our experimental results on the impact of edge removal, generated through the node2vec embedding method with dimension $16$, are reported in Fig. 13. Our experimental results using node2vec embedding method with dimension $32$ are collected in Fig. 23 (see Appendix A). The heatmaps in Fig. 14 illustrate that node2vec proves to be a robust method to produce community detection outcomes with high ECS, particularly for strongly clustered networks. In Fig. 14(b) for weakly clustered networks (i.e., $\mu=0.5$), to get high ECS, in general, lower embedding dimensions produce more robust community partition for networks with small sizes; higher embedding dimensions produce more robust community partition for networks with larger sizes.
@@
Figure 13: Mean ECS using node2vec method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. Random node selection is applied in (a) and (b), while (c) and (d) use targeted selection. (a) and (c) deal with $1,000$-node networks, and (b) and (d) with $10,000$-node networks.
Figure 14: Heatmaps with node2vec embedding method. For given embedding dimension and LFR network with given size, the ECS similarity score between ground-truth community structure and detected community structure is recorded.
III.2.4 Summary of random walk methods
We derive the following insights for random walk methods. In the cases of LINE and node2vec, we consistently observe a decrease in ECS as we iteratively select additional nodes and remove their adjacent edges, which aligns with our intuitive expectations. However, DeepWalk exhibits some unexpected increases in ECS. Moreover, for smaller values of the mixing parameter $\mu$, the corresponding curve tends to exhibit higher values. When employing targeted node selection, the ECS experiences a more significant and rapid decline contrasted with random node selection. For example, for LFR network with $1,000$ nodes and $\mu=0.5$, when perturbing the network at $50\%$ and using node2vec with $16$-dimensional embedding, we notice that targeted node selection decreases ECS from $0.43$ to $0.13$ (a decrease of $69.8\%$, see Fig. 13(c)), but ECS decreases from $0.43$ to $0.25$ for random node selection (a decrease of $41.9\%$, see Fig. 13(a)).
From our discussion of each method, we observe that DeepWalk may exhibit an unexpected increase while LINE experiences unexpected oscillations. In contrast, node2vec is free from these issues and consistently achieves the highest ECS similarity scores. When considering the application of random walk embeddings for community detection, node2vec consistently proves to be a more robust choice over DeepWalk and LINE, across networks of varying sizes, the strength of the initial community partition, and perturbation type.
III.3 Comparison of graph embedding methods
In preceding sections, we utilized heatmaps to assess the performance of graph embedding methods for various network sizes and embedding dimensions. To compare community detection outcomes given by various embedding techniques, we conduct cross-comparisons between six data clustering methods. Due to the high computational cost associated with the LINE embedding method, we exclude it from our comparative analysis. The corresponding heatmap for cross-comparison is depicted in Fig. 15. Specifically, Fig. 15(a) comprises a $10\times 5$ matrix, totaling $50$ entries, which we interpret as a $50$-dimensional vector for analysis. Thus, for $\mu=0.01$ or $\mu=0.5$, we obtained a $50$-dimensional vector for each data clustering method based on LE, LLE, HOPE, M-NMF, DeepWalk, and node2vec. These six vectors (given by six data clustering methods) underwent cross-comparison by calculating cosine similarity. In other words, heatmap cells in Fig. 15 reveal the similarity between data clustering results obtained by different methods. Notably, we observe that LE and LLE yield similar data clustering results. However, the results obtained by HOPE were not as closely aligned with those by LE and LLE, although HOPE is also in the family of matrix factorization methods. DeepWalk and node2vec, as two methods in the family of random walk methods, also exhibit similar results. Remarkably, M-NMF yield results significantly divergent from those produced by other methods.
Figure 15: Cross comparison between data clustering methods.
III.4 Results on real-world networks
In this section, we present the results of our experiments conducted on two real-world networks with labeled communities, i.e., the email-EU-core network and the AS network. The email-EU-core network is constructed using email data sourced from a prominent European research institution, comprising $986$ nodes and $16,687$ edges [63]. The AS network, derived from the AS Internet topology data collected in June 2009 by the Archipelago active measurement infrastructure, represents the inter-domain Internet topology with Autonomous Systems (ASes) as nodes and AS peerings as links, forming an AS-level topology graph. This dataset comprises $23,748$ nodes and $58,414$ edges. We select them because they closely resemble power-law degree and community size distributions, as the synthetic networks counterpart (i.e., LFR benchmark graphs). To confirm the alignment of these real-world networks with the properties of LFR networks, we conducted statistical tests using the Python package powerlaw based on the statistical test derived in the work by Clauset et al [66]. The results indicate that both the email-EU-core network and the AS network exhibit power-law characteristics [67]. Specifically, in the email-EU-core network, the degree distribution adheres to a power-law distribution and the community size sequence follows an exponential distribution; while in the AS network, both degree and community size distributions conform to power laws.
As discussed in Section II.2, in experiments involving LFR networks, we ensure that the removal of edges from each set of nodes does not result in the disconnection of the remaining network — a prerequisite for graph embedding methods. However, such a stringent requirement is often impractical for real-world networks. For instance, randomly selecting $5\%$ of nodes in the email-EU-core network and deleting their adjacent edges would typically leave the remaining portion of the network disconnected. To address this issue in our experiments with real-world networks, we adopt an alternative approach. In particular, we randomly introduce edges between connected components to restore connectivity. Subsequently, we proceed with the embedding of the reconnected network.
We report experimental results using different methods with $16$-dimensional embedding on these two real-world networks in Fig. 16. Comparing the two rows of Fig. 16, we can see that when employing targeted node selection, the ECS experiences a more significant decrease than the case of random node selection. For example, for the email-EU-core network, when perturbing it at $50\%$ and using node2vec with $16$-dimensional embedding, we notice that targeted node selection decreases ECS from $0.41$ to $0.31$ (a decrease of $24.3\%$, see Fig. 16(c)), but ECS is only decreased by around $0.02$ by random node selection (Fig. 16(a)). Furthermore, LLE and node2vec proved to be the superior choices among all seven embedding methods. In fact, curves corresponding to these two methods in Fig. 16 are higher in position than other curves in general. The experimental results using different methods with $32$-dimensional embedding on the two real-world networks are depicted in Fig. 24.

### v3

@@ III.4 Community Similarity
Network Generation, Sec. III.1Perturbation, Sec. III.2Data Clustering, Sec. III.3Similarity, Sec. III.4
@@
IV Results and Discussion
In this section, we present the results of our experiments. Recall that Sec. III.3 details the embedding methods we use here. In the following subsections, we use $\mathbf{A}$ to denote the adjacency matrix of a graph. We report results based on the family of graph embedding methods (i.e., matrix-based family in Sec. IV.1 and random walk family in Sec. IV.2).
In the subsequent line figures, unless otherwise stated (in Figs. 17, 18 and 14), subfigures in the top row, i.e., (a) and (b), correspond to random node selection. More specifically, subfigure (a) corresponds to experimental results on small networks with 1,0001,0001 , 000 nodes and subfigure (b) depicts experimental results on large networks with 10,00010,00010 , 000 nodes. Similarly, subfigures in the bottom row, i.e., (c) and (d), correspond to targeted node selection, for small and large networks respectively.
We present line figures using 161616-dimensional and 323232-dimensional embedding. We focus only in these two dimensions as smaller embedding dimensions seem to be enough to provide an accurate low-dimensional vector representation of networks [15, 16]. Recall that a higher ECS value indicates a greater similarity between network community partitions obtained from the LFR benchmark graphs and the embeddings. Thus, when analyzing line figures, higher ECS values and therefore higher positions of line curves indicate a more robust graph embedding method for community detection.
Table 2 shows the experimental parameters employed by the embedding methods, excluding embedding dimension, $d$. Notably, LE and LLE methods only require the embedding dimension $d$. For the remaining parameters, we utilized the default parameters as specified in Tandon et al. (2021) [14]; they show that default values of parameters for the embedding methods lead to comparable performance as traditional community detection methods, and that the optimized parameters do not exhibit a significant improvement over default parameters.
@@ IV Results and Discussion
Random Walk LINE batch size: 819281928192, epochs: 100100100
@@
Finally, Sec. IV.3 presents experimental results to evaluate the robustness of our community detection methods on two real-world networks with labeled communities: the email-EU-core network [77] and the AS network [7]. These networks exhibit similar characteristics to synthetic counterparts, specifically power-law degree and community size distributions as observed in LFR benchmark graphs. Our experimental methodology and insights for real-world networks mirror findings for synthetic counterparts.
IV.1 Matrix Factorization Methods
@@ IV.1.1 Laplacian Eigenmap (LE)
$\mathbf{D}^{-1}(\mathbf{D}-\mathbf{A}).$
Our experimental results on the impact of edge removal, generated through the LE embedding method with dimension 161616, are reported in Fig. 2. Our experimental results using LE embedding method with dimension 323232 are collected in Fig. 3.
Figure 2: Mean ECS using LE method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. Random node selection is used in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) are for networks with 1,0001,0001 , 000 nodes, and (b) and (d) are for networks with 10,00010,00010 , 000 nodes.
Figure 3: Mean ECS using LE method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. Random node selection is used in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) are for networks with 1,0001,0001 , 000 nodes, and (b) and (d) are for networks with 10,00010,00010 , 000 nodes.
[…]
IV.1.2 Locally Linear Embedding (LLE)
@@
Our experimental results on the impact of edge removal, generated through the LLE embedding method with dimensions 161616 and 323232, are reported in Fig. 4 and Fig. 5, respectively.
Figure 4: Mean ECS using LLE method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. (a) and (b) apply random node selection, whereas (c) and (d) use targeted selection. Networks in (a) and (c) have 1,0001,0001 , 000 nodes, and those in (b) and (d) have 10,00010,00010 , 000 nodes.
Figure 5: Mean ECS using LLE method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. (a) and (b) apply random node selection, whereas (c) and (d) use targeted selection. Networks in (a) and (c) have 1,0001,0001 , 000 nodes, and those in (b) and (d) have 10,00010,00010 , 000 nodes.
The insights drawn from Figs. 4 and 5 for LLE align closely with the insights derived for LE. In fact, the curves in Fig. 4 are slightly higher in position than those in Fig. 2, indicating that LLE outperforms LE. For example, the data points on the green curve (corresponding to $\mu=0.4$) from LLE in Fig. 4(a) have higher ECS scores over LE in Fig. 2(a) for a given percentage of selected nodes.
IV.1.3 Higher-Order Preserving Embedding (HOPE)
@@
Our experimental results on the impact of edge removal, generated through the HOPE embedding method with dimensions 161616 and 323232, are reported in Fig. 6 and Fig. 7, respectively.
Figure 6: Mean ECS using HOPE method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. While (a) and (b) use random node selection, (c) and (d) rely on targeted selection. Networks in (a) and (c) contain 1,0001,0001 , 000 nodes, while those in (b) and (d) consist of 10,00010,00010 , 000 nodes.
Figure 7: Mean ECS using HOPE method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. While (a) and (b) use random node selection, (c) and (d) rely on targeted selection. Networks in (a) and (c) contain 1,0001,0001 , 000 nodes, while those in (b) and (d) consist of 10,00010,00010 , 000 nodes.
For the HOPE embedding method, our results suggest that using a 32-dimensional embedding (i.e., Fig. 7) brings higher ECS scores than a 16-dimensional one (i.e., Fig. 6). For example, the data points on the green curve (corresponding to $\mu=0.4$) in Fig. 7(a) have higher ECS scores than those data points on the green curve (corresponding to $\mu=0.4$) in Fig. 6(a) at respective percentage of selected nodes. We can also see that HOPE has higher ECS scores with smaller graphs.
IV.1.4 Modularized Non-Negative Matrix Factorization (M-NMF)
@@
Our experimental results on the impact of edge removal, generated through the M-NMF embedding method with dimensions 161616 and 323232, are reported in Fig. 8 and Fig. 9, respectively.
Figure 8: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. (a) and (b) implement random node selection, while (c) and (d) apply targeted selection; (a) and (c) correspond to networks with 1,0001,0001 , 000 nodes, and (b) and (d) to networks with 10,00010,00010 , 000 nodes.
Figure 9: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. (a) and (b) implement random node selection, while (c) and (d) apply targeted selection; (a) and (c) correspond to networks with 1,0001,0001 , 000 nodes, and (b) and (d) to networks with 10,00010,00010 , 000 nodes.
From Figs. 8 and 9, we can see that M-NMF produces higher ECS scores with smaller graphs.
@@
Larger Network (larger $N$) LE with a higher embedding dimension LE or LLE with a lower embedding dimension
IV.1.5 Summary of Matrix Factorization Methods
[…]
For smaller real-world graphs containing a few thousand vertices, to obtain more robust partitions, the LLE method is suggested with a modest embedding dimension, typically on the order of tens, such as 16 dimensions; Alternatively, when the community partition is stronger, the HOPE embedding method with a higher embedding dimension produces comparably robust partitions. In the case of larger real-world graphs with a weaker community partition, to obtain more robust community partitions, it is suggested to use the LE embedding method with a moderately higher embedding dimension, such as 323232 dimensions as utilized in our experiments (see Fig. 3). When dealing with larger real-world graphs characterized by a stronger community structure, the LLE or LE embedding method with a lower embedding dimension produces more robust community partitions. In summary, LLE is the most robust method among matrix factorization methods. Table 3 summarizes our findings for matrix factorization methods.
IV.2 Random Walk Methods
Random walk methods learn network embeddings of graph nodes by modeling a stream of short random walks. We chose DeepWalk [31], LINE [32], and node2vec [33] as widely used methods of this family.
IV.2.1 DeepWalk
@@
Our experimental results on the impact of edge removal, generated through the DeepWalk embedding method with dimension 161616, are reported in Fig. 10. Our experimental results using DeepWalk embedding method with dimension 323232 are collected in Fig. 11.
Figure 10: Mean ECS using DeepWalk method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings.Random selection is employed in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) represent networks with 1,0001,0001 , 000 nodes, and (b) and (d) represent those with 10,00010,00010 , 000 nodes.
Figure 11: Mean ECS using DeepWalk method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. Random selection is employed in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) represent networks with 1,0001,0001 , 000 nodes, and (b) and (d) represent those with 10,00010,00010 , 000 nodes.
From Figs. 10 and 11 for DeepWalk, we can observe that networks with stronger initial partitions (i.e., lower $\mu$ values) experience less perturbation impact and targeted node selection has a significant influence on clusters. Moreover, from the right column of Fig. 11, we see that some curves exhibit unexpected increase as we progressively selected more nodes.
IV.2.2 Large-Scale Information Network Embedding (LINE)
@@
where $E$ is the edge set and $\mathbf{x}_{i}$ denotes the embedding point of node $i$.
The optimization process of the LINE method utilizes stochastic gradient descent, which is enhanced by an edge-sampling treatment, as detailed in [32]. Specifically, the utilization of the LINE embedding method necessitates the presence of GPU for computation. In our study, the machine learning model was trained using a batch size of 819281928192 over 100100100 epochs, facilitating iterative updates to the embedding vector across the entire training dataset.
Our experimental results on the impact of edge removal, generated through the LINE embedding method with dimensions 161616 and 323232, are reported in Fig. 12 and Fig. 13, respectively.
Figure 12: Mean ECS using LINE method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. (a) and (b) follow random node selection, whereas (c) and (d) adopt targeted node selection. Networks with 1,0001,0001 , 000 nodes correspond to (a) and (c), and those with 10,00010,00010 , 000 nodes correspond to (b) and (d).
Figure 13: Mean ECS using LINE method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. (a) and (b) follow random node selection, whereas (c) and (d) adopt targeted node selection. Networks with 1,0001,0001 , 000 nodes correspond to (a) and (c), and those with 10,00010,00010 , 000 nodes correspond to (b) and (d).
From Figs. 12 and 13. the LINE embedding method produces higher ECS scores when applied to larger networks. Our analysis of networks with 1,000 nodes in the first column of Fig. 12 reveals that curves exhibit a steep decline followed by a subsequent increase, but we notice that increasing the embedding dimensions in LINE is expected to mitigate the sharpness of this decline and increase, rendering the pattern less pronounced, as indicated in Fig. 14. Based on the aforementioned observation, LINE becomes a viable option for network clustering when dealing with relatively larger networks and the computational resources, particularly GPUs, are available.
Figure 14: Mean ECS using LINE method over the percentage of selected nodes uniformly at random on LFR benchmark graphs with 1,0001,0001 , 000 nodes. The first plot displays results using 128128128-dimensional embeddings; the second plot displays results using 256256256-dimensional embeddings.
IV.2.3 node2vec
@@
Our experimental results on the impact of edge removal, generated through the node2vec embedding method with dimensions 161616 and 323232, are reported in Fig. 15 and Fig. 16, respectively.
@@
Figure 15: Mean ECS using node2vec method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. Random node selection is applied in (a) and (b), while (c) and (d) use targeted selection. (a) and (c) deal with 1,0001,0001 , 000-node networks, and (b) and (d) with 10,00010,00010 , 000-node networks.